Simplify: ∛64/∜256


Answer by rothauserc(4718)

64^(1/3) * 256^(1/4) = 4 * 4 = 16
:
Note x^(1/3) says find the cube root of x, x^(1/4) means find the 4th root of x and x^(1/2) means find the square root of x
:
therefore,
:
4 * 4 * 4 = 64, so 64^(1/3) = 4
4 * 4 * 4 * 4 = 256, so 256^(1/4) = 4
: